find the equation of the tangent to the curve


exy=x+(In y)2  at the point (0,e)

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

by implicit differentiation,(yexy+xexy)dy/dx =1+(2lny)/y,at (0,e),edy/dx=1+(2/e),dy/dx=(1+(2/e))/e, the remaining part of the solution is straightforward.